OpenStudy (anonymous):

square root of 368 simplified?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt{368}=\sqrt{2^4\cdot23}=2^2\sqrt{23}=4\sqrt{23}\]